# DeepDISC-photoz: Deep Learning-Based Photometric Redshift Estimation for Rubin LSST

### v1

@@ 4.2 Observational Systematics
We look at the sensitivity to imaging systematics in Figures 8, 9, 10, namely the Milky Way dust reddening E(B-V), blendedness, and point spread function (PSF) full width at half-max (FWHM) in the i-band.
Milky Way dust in the line of sight of a source will scatter and absorb light and thermally re-emit in the infrared. Observationally, this will appear as a dimming and reddening (Draine, 2003). For catalog-based photo-z codes that use measured fluxes or magnitudes, it is important to correct for this effect by using the measured reddening E(B-V) and extinction law for each photometric filter. The DC2 area was chosen to be at high Galactic latitude in a relatively uniform, low-dust region, which results in the small maximum E(B-V) and apparent bimodality in the distribution. We see a slightly positive trend in bias vs E(B-V) for DeepDISC and an uptick in outlier fraction at the highest E(B-V) values. However, the bias remains minimal and below the LSST science requirements error budget. No strong trend is seen in scatter. FZB also appears largely insensitive to E(B-V), but BPZ metrics tend to increase as E(B-V) does.
@@
where $flux_{\rm child}$ is the deblended child flux and $flux_{\rm parent}$ is the total flux in the footprint. Isolated sources thus have a blendedness of 0, and a blendedness above 0.2 can be considered a significant blend. For details of the blendedness metric, see Bosch et al. (2018). The effect of blending on all codes is shown in Figure 9. Up to a blendedness of around 0.5, BPZ bias increases, FZB bias slightly increases, and DeepDISC bias remains roughly constant. However, beyond a blendedness of 0.5, BPZ and FZB bias start to dramatically decrease, while DeepDISC bias only slightly decreases. For all codes, photo-z scatter increases as a function of blendedness, although DeepDISC is the least sensitive to this systematic. It appears that the features extracted by DeepDISC are more robust to blending than the deblended cModel magnitudes produced by the LSST science pipelines deblender. This may be due to the fact that all pixel-level information is considered in DeepDISC feature extraction, which provides more information regarding source light profiles and SEDs than aperture magnitude measurements. We discuss this more in Appendix A.
We examine any systematics trends due to i-band PSF FWHM in Figure 10. The PSF characterizes the distortion of an observed image due to atmospheric effects such as refraction. Since DeepDISC uses image data and thus is encoding pixel-level features and distortions, we examine if there is any dependence in performance with PSF FWHM. No clear trends can be seen in Figure 10, except a potential dip in outlier fraction at high values. Overall, DeepDISC shows little sensitivity to these systematics, comparable to the catalog-based methods which incorporate systematics mitigation during pre-processing, e.g., deblending, or post-processing, e.g., correcting for reddening.
@@ 4.3 Dependence on Image Depth
Figure 12: Redshift point estimates for Year 1 (left) and Year 5 (right) data. The mode is used to produce a single point estimate for each pdf. The dashed line is along $z_{\rm true}=z_{\rm phot}$, and the solid lines define the $3\sigma_{\rm IQR}$ outlier boundary. Color corresponds to the number density of objects in each bin. Bias, scatter and outlier fraction (Equations 3, 4 and 5, respectively) are shown in the legends. The comparison is made on the common set of objects detected by both models. Year 5 data greatly improves the estimates overall, and notably improves the high-z regime.
@@ 4.4 Model Scalability
An investigation of the scaling laws governing our DeepDISC model is shown in Figures 13 and 14. We train using the Year 5 data and show the effects of independently increasing the model and training set size, quantified by the photo-z metrics used thus far. Overall, there does not appear to be a strong scaling law for either model size or training set size, in terms of both point estimate and PDF metrics. This is mostly consistent with Walmsley et al. (2024), as they find that while models pre-trained on ImageNet modestly benefit from increased model size, the largest gains were realized with pre-training in-domain on astronomical images. Our baseline model appears to have reached a "saturation" of possible gains with increased size or data, indicating other avenues to improve performance may be more fruitful. A full investigation of how DeepDISC scales in different pre-training contexts merits is left to future work. Here, we are testing whether a simple change to the training set or model size yields noticeable benfits.
@@ 5 Conclusions
In order to conduct a fair comparison of the codes, we supply the other codes the object catalogs, which have been reduced from the input images. This ensures all methods are trained and applied to data that accounts for imaging systematics such as blending. We find that DeepDISC outperforms the catalog-based codes in almost all metrics on the test set, including metrics designed to evaluate the quality of point estimates and uncertainty predictions. We find that DeepDISC photo-z tends to under-predict high redshifts and does not produce predictions above $z\sim 2.5$, likely due to the low representation of these objects in the training. However, compared to FlexZBoost, which is also a machine learning method sensitive to the training distribution, we find DeepDISC produces similar (although better) bias, scatter, and outlier fraction using the mode of the predicted PDFs.
We compare the quality of the predicted redshift PDFs of all codes by examining the PIT histograms and probabilistic metrics. DeepDISC produces the empirical PIT histogram closest to the ideal theoretical uniform distribution, shown in the visualization and statistical measurements. We find that we are able to recover over 51 percent of photo-z outliers (i.e., they are no longer considered outliers) if $z_{\rm phot}$ is taken be the secondary peak of the PDF, indicating that meaningful degeneracies are learned by the network.
@@
In addition to the tests above, we examine possible scaling laws that could govern our model. The backbone of the network is a MViTv2 network, a vision transformer which is known for being data hungry (Zhai et al., 2022). We investigate if our models (pretrained on ImageNet) have notable performance gains when the data set size or model size is increased. We find no strong relation for any metric, point or probabilistic, as a function of model size or training set size. Scaling studies often find power law dependencies, so it is possible we are already in the saturated regime as we do not see any notable gains. Walmsley et al. (2024) found that while pre-training on ImageNet data can improve downstream galaxy morphology tasks and does scale with data set and model size, most performance gains come from additional pre-training on galaxy images. Investigation into this effect for DeepDISC will be a promising avenue of future work, especially as pre-training using galaxy images has been shown to improve the detection of galactic features through a Faster-RCNN network (Popp et al., 2024). Pre-training on simulated images (such as those used in this study) may help with training the model on real data, which is incomplete in redshift and color-magnitude space due to spectroscopic selection effects. Training sample augmentation has been shown to mitigate this bias and it would be worthwhile to explore whether DeepDISC would benefit from this method. Moskowitz et al. (2024) augmented a photometric catalog to produce a new training set by using a normalizing flow, and such a catalog could be used to simulate new images to produce an augmented training set for DeepDISC. DeepDISC has shown promising performance on simulated LSST data, which we aim to continue to improve and vet as we inevitably approach Rubin first light.
@@ Acknowledgements
This paper has undergone an internal review by the LSST DESC, and we thank the internal reviewers, Eric Gawiser and Ismael Mendoza, for thoughtful feedback that improved the paper. We thank Athol Kemball, Matias Carrasco Kind, and Yuxiong Wang for illuminating discussions. We thank Shirui Luo, Dawei Mu, and Volodymyr Kindratenko at the National Center for Supercomputing Applications (NCSA) for their assistance with the GPU cluster used in this work. G.M., X.L., Y.E. and Y.L. acknowledge support from the NCSA Faculty Fellowship, NCSA Student Pushing INnovation internship program, LSST-DA through grant 2023-SFF-LFI-03-Liu, NSF grant AST-2308174, and NASA grant 80NSSC24K0219. G.M. thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, and the Moore Foundation; his participation in the program has benefited this work. This work was conducted as part of a LINCC Frameworks Incubator. LINCC Frameworks is supported by Schmidt Sciences. AIM, TZ, DO, OL, DB, JK, and MD are supported by Schmidt Sciences. C.J.B. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303803. This material is based upon work supported by the National Science Foundation under Award No. 2303803. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. This work utilizes resources supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign.

### v2

@@ 4.2 Observational Systematics
Due to the many exposures that will be collected by LSST over its observing run, the image coaddition process will result in an averaging over several systematics. However, it is still important to characterize the performance of our photo-z estimator as a function of imaging systematics to understand weaknesses and biases that may propagate to downstream measurements. In this work, we have only explicitly accounted for Milky Way dust reddening by adding the E(B-V) values as extra input to the redshift ROI head of DeepDISC. We examine photo-z performance as a function of E(B-V) as well as point spread function (PSF) variations and blending in Figures 8, 9 and 10.
Milky Way dust in the line of sight of a source will scatter and absorb light and thermally re-emit in the infrared. Observationally, this will appear as a dimming and reddening (Draine, 2003). For catalog-based photo-z codes that use measured fluxes or magnitudes, it is important to correct for this effect by using the measured reddening E(B-V) and extinction law for each photometric filter. The DC2 area was chosen to be at high Galactic latitude in a relatively uniform, low-dust region, which results in the small maximum E(B-V) and apparent bimodality in the distribution in Figure 8. We see a slightly positive trend in bias vs E(B-V) for DeepDISC and an uptick in outlier fraction at the highest E(B-V) values. However, the bias remains minimal and below the LSST science requirements error budget. No strong trend is seen in scatter. FZB and BPZ also appear largely insensitive to E(B-V), which is expected due to the aforementioned dust correction.
@@
where $flux_{\rm child}$ is the deblended child flux and $flux_{\rm parent}$ is the total flux in the footprint. Isolated sources thus have a blendedness of 0, and a blendedness above 0.2 can be considered a significant blend. For details of the blendedness metric, see Bosch et al. (2018). The effect of blending on all codes is shown in Figure 9. Up to a blendedness of around 0.5, BPZ bias increases, FZB bias slightly increases, and DeepDISC bias slightly decreases. However, beyond a blendedness of 0.5, BPZ and FZB bias start to dramatically decrease, while DeepDISC bias only slightly decreases. For all codes, photo-z scatter increases as a function of blendedness, although DeepDISC is the least sensitive to this systematic. It appears that the features extracted by DeepDISC are more robust to blending than the deblended cModel magnitudes produced by the LSST science pipelines deblender. This may be due to the fact that all pixel-level information is considered in DeepDISC feature extraction, which provides more information regarding source light profiles and SEDs than aperture magnitude measurements. We discuss this more in Section 5 and Appendix A. Additionally, raising the positive IOU threshold discussed in Section 3 could prevent more detection/truth mismatches during training, which could further improve the robustness to this systematic.
We examine any systematics trends due to i-band PSF FWHM in Figure 10. The PSF characterizes the distortion of an observed image due to atmospheric effects such as refraction. Traditional image analysis pipelines often explicitly use the PSF as input, potentially giving them an advantage over neural network approaches that do not, as is the case with DeepDISC. Since DeepDISC uses image data and thus is encoding pixel-level features and distortions, we examine if there is any dependence in performance with PSF FWHM. No clear trends can be seen in Figure 10, except a potential dip in outlier fraction at high values. This is encouraging, in that it shows that training with a variety of images and using a “data-driven" approach is sufficient to account for (small) PSF variations rather than explicitly inputting this information into the network.
Overall, DeepDISC shows little sensitivity to these systematics, comparable to the catalog-based methods which incorporate systematics mitigation during pre-processing, e.g., deblending, or post-processing, e.g., correcting for reddening. We find that DeepDISC in particular is more robust to blending than the catalog-based codes, with significantly less photo-z scatter at high levels of blending. In future applications with real data, a variety of images taken from a wide footprint should be used in order to help the network marginalize over varying systematics. Understanding the photo-z uncertainties due to imaging systematics will be an important endeavor for downstream science cases.
@@ 4.3 Dependence on Image Depth
Figure 12: Redshift point estimates for Year 1 (left) and Year 5 (right) data. The mode is used to produce a single point estimate for each pdf. The dashed line is along $z_{\rm true}=z_{\rm phot}$, and the solid lines define the $3\sigma_{\rm IQR}$ outlier boundary. Color corresponds to the number density of objects in each bin. Bias, scatter and outlier fraction (Equations 4, 5 and 6, respectively) are shown in the legends. The comparison is made on the common set of objects detected by both models. Year 5 data greatly improves the estimates overall, and notably improves the high-z regime.
@@ 4.4 Model Scalability
An investigation of the scaling laws governing our DeepDISC model is shown in Figures 13 and 14. We train using the Year 5 data and show the effects of independently increasing the model and training set size, quantified by the photo-z metrics used thus far. Overall, there does not appear to be a strong scaling law for either model size or training set size, in terms of both point estimate and PDF metrics. This is mostly consistent with Walmsley et al. (2024), as they find that while models pre-trained on ImageNet modestly benefit from increased model size, the largest gains were realized with pre-training in-domain on astronomical images. Our baseline model appears to have reached a "saturation" of possible gains with increased size or data, indicating other avenues to improve performance may be more fruitful. A full investigation of how DeepDISC scales in different pre-training contexts merits is left to future work. Here, we are testing whether a simple change to the training set or model size yields noticeable benefits.
5 Discussion
Figure 15: After DeepDISC inference is ran, a small sample is set aside and divided into training and test samples for FZB. We compare the mode point estimate results on this small test sample. The dashed line is along $z_{\rm true}=z_{\rm phot}$, and the solid lines define the $3\sigma_{\rm IQR}$ outlier boundary. DeepDISC is shown on the left, and FZB trained on the small training sample photometry/DeepDISC features is shown in the middle and right, respectively. Training with DeepDISC features leads to reduced scatter and outlier fractions in the test sample. DeepDISC also produces lower photo-z scatter than FZB trained on the DeepDISC-selected sample.
Figure 16: PIT histograms and QQ plots of the PDF ensembles produced by DeepDISC (left), and FZB trained on photometry (middle) vs DeepDISC features (right). PIT histograms are shown in grey and correspond to the right y-axis. The black horizontal line represents a uniform PIT histogram. The red line plots quantiles of the ensemble PIT values compared to quantiles of an ideal flat distribution and corresponds to the left y-axis.
DeepDISC photo-z is designed as an end-to-end pipeline for downstream photo-z measurements. Like other computer vision models, DeepDISC skips the need for photometry as an intermediate step and directly produces a photo-z catalog. The features that the DeepDISC backbone learns during training do not have obvious physical meaning, but do capture relevant information. In the previous sections, we examined DeepDISC performance as a function of object properties and imaging systematics to quantify and understand some of the performance gains relative to traditional catalog-level photo-z algorithms. DeepDISC appears to be able to handle cases in which traditional algorithms struggle, such as highly blended objects and regions of redshift space in which strong spectral features are not covered by the optical LSST filters. This performance gain is likely due to the extra information gathered from pixel values and extracted by the backbone network, as opposed to traditional aperture photometry.
We test this hypothesis by decoupling the photo-z estimation from the feature extraction. To do this, we compare an FZB model trained on photometry with an FZB model trained on the corresponding features produced by a trained DeepDISC model. The feature tensor output by the backbone is large (256x7x7 for each object) so we first flatten it and run principle component analysis (PCA) to reduce the dimensionality of the features to 12 values per object, matching the size of the photometric catalog (magnitudes + errors for 6 filters). Some spatial information may be lost due to the flattening, so the results here can be considered a lower bound on the relative performance. We run the trained DeepDISC model on a small subset of images from the main test set. We perform the same catalog-matching and dust reddening corrections described in Section 3.2 and Section 4 to produce a matched catalog of photometry and corresponding DeepDISC features. We then divide this catalog into a training and test set for FZB. We train FZB using the same hyperparameters described in Section 3.2, varying only whether the model uses photometry or DeepDISC features.
The test set point estimates for DeepDISC and both models are shown in Figure 15. The FZB model trained with DeepDISC features produces less scatter in its point estimates compared to the model trained with photometry (middle versus right panels). PIT histograms for the models are shown in Figure 16. Differences are small, but the model trained with DeepDISC features produces a slightly flatter histogram, and less outliers at the edge of the distribution. Overall, we find that the backbone network is able to extract more information relative to general photo-z estimation than colors/magnitudes derived from aperture photometry.
5.1 Sample Selection
In the analysis in Section 4, the BPZ and FZB methods were trained on catalog photometry matched to the DC2 truth catalog that DeepDISC was trained on. However, the test set selection is determined by DeepDISC inference matched with the DC2 object catalog, potentially privileging our method as it determines the sample selection that the other codes use for inference. I.e., the selection effects may make the training data non-ideal for FZB and BPZ in the previous analysis. This effect is accounted for in the analysis above, as training and test data for FZB is taken from the DeepDISC selection catalog. Comparing the left and middle panels of Figures 15 and 16 shows the relative results when this selection is accounted for. DeepDISC still produces a lower scatter than FZB, but its bias is slightly higher (although still within LSST science requirements). The PIT histogram appears slightly flatter for DeepDISC, with less outliers at the edges. We further discuss the relevance of sample selection in the context of weak lensing below.
5.2 Weak Lensing
For studies of weak lensing that require tomographic binning of redshifts to produce $n(z)$ distributions, characterizing major sources of bias in the weak lensing measurement is necessary so that they may be accounted for and mitigated (Mandelbaum, 2018). Particularly relevant for weak lensing with DeepDISC photo-zs are selection biases due to source detection and photo-z estimation. Source detection may implicitly depend upon galaxy shape through its correlation with lensing. The photo-z estimation may also be dependent on galaxy shape due to the incorporation of morphological information. A rigorous investigation of these effects and a full weak lensing analysis is deserving of its own work and outside the current scope here. However, selection biases have been well-studied in the literature, so there are paths forward.
One way to account for selection effects is through self-calibration, commonly done with the metacalibration algorithm (Huff & Mandelbaum, 2017; Sheldon & Huff, 2017). This process involves artificially applying a shear to real galaxy images and measuring the response of some measurement to the shear. Selection biases can then be quantified and removed from subsequent weak lensing analysis. metacalibration has been applied to galaxy shape estimators (Gatti et al., 2021), photo-z estimators (Troxel et al., 2018), and object detection algorithms (Sheldon et al., 2020). Future work will adapt DeepDISC to simultaneously estimate redshift and galaxy shapes, providing a single method to detect, deblend, measure and calibrate selection biases.
5.3 Limitations
The simulations used in this work include many realistic effects, both due to physical phenomena and instrumental design. While they encapsulate an enormous effort to represent the universe as close to the “truth” as possible, there are still fundamental limitations in our experiment and thus potential differences in how the model will perform on real data compared to the results here. As described in Section 2, the morphological model for objects in the DC2 simulations is a simple bulge+disk+knot model. The size-luminosity relation for disk and bulge components is determined from fitting a half-light radius to the functional form of Zhang & Yang (2019), such that parameters for SDSS galaxies are recovered at $z=0$ and sizes decrease with redshift following a sigmoid function. This relation, along with the simple morphological profiles, cannot capture the true diversity of galaxy shapes and sizes that will be observed by LSST. However, higher fidelity simulations with generative models or real morphological profiles may help to bridge the simulation-reality gap. We are also encouraged by preliminary results on real JWST images (Merz et al., in prep).
In this work, our training sample is completely representative of our test sample, which is a best case scenario for machine learning methods like DeepDISC and FZB. It is worthwhile to note that in these ideal conditions, DeepDISC outperforms FZB, but the metrics can in some sense be thought of as upper bounds on the performance. Moskowitz et al. (2024) show that underrepresented training samples can severely bias FZB photo-z estimation, and since DeepDISC is also a machine-learning estimator, is it likely also susceptible to this effect. The solution proposed by Moskowitz et al. (2024) is to use data augmentation to help the model learn in underrepresented regimes. For image-based methods such as DeepDISC, simulated images could be generated from augmented catalogs, or the DC2 images themselves could be used to give the networks a more complete picture of the color-redshift space.
Figure 17: Mode point estimates of the DeepDISC model trained with 5x the amount of training data. The network is able to better estimate photo-zs above $z_{\textrm{phot}}=2.5$.
One of the main limitations of DeepDISC photo-z is evident in the lack of PDF modes above $z=2.5$. While all codes struggle in this high-z regime, it is particularly stark in the DeepDISC scatter plot in Figure 4. Photo-z estimation in this regime is fundamentally a challenge, as the low sample size of objects limits the network’s ability to learn relevant features. One potential avenue to improve DeepDISC estimates in the high-z regime is through increasing the amount of high-z data in the training sample. While we investigated the scaling laws in Section 4.4 and found no strong improvement of point estimate metrics with increased training set size, those metrics were calculated over the whole sample and did not capture the behavior of the networks in different regimes. We find that the models trained with 2x and 5x training data are able to provide mode point estimates beyond the $z=2.5$ cutoff (see Figure 17). This test increased the number of high-z objects in the training set, but preserved the overall $n(z)$. In future work, targeted data augmentation like that of Moskowitz et al. (2024) could be implemented in order to specifically increase $n(z>2.5)$. The models are very sensitive to regimes with low representation, so added data does make a difference. Different pre-training schemes and architectures may also lead to improved high-z performance and will be explored in follow-up work.
@@ 6 Conclusions
In order to conduct a fair comparison of the codes, we supply the other codes the object catalogs, which have been reduced from the input images. This ensures all methods are trained and applied to data that accounts for imaging systematics such as blending. We find that DeepDISC outperforms the catalog-based codes in almost all metrics on the test set, including metrics designed to evaluate the quality of point estimates and uncertainty predictions. We find that DeepDISC photo-z tends to under-predict high redshifts and does not produce mode point estimates above $z\sim 2.5$, likely due to the low representation of these objects in the training. However, this may be ameliorated with additional high-z data augmentation.
We compare the quality of the predicted redshift PDFs of all codes by examining the PIT histograms and probabilistic metrics. DeepDISC produces the empirical PIT histogram closest to the ideal theoretical uniform distribution, shown in the visualization and statistical measurements. We find that we are able to recover a majority of photo-z outliers (i.e., they are no longer considered outliers) if $z_{\rm phot}$ is taken be the secondary peak of the PDF, indicating that meaningful degeneracies are learned by the network.
@@
In addition to the tests above, we examine possible scaling laws that could govern our model. The backbone of the network is a MViTv2 network, a vision transformer which is known for being data hungry (Zhai et al., 2022). We investigate if our models (pre-trained on ImageNet) have notable performance gains when the data set size or model size is increased. We find no strong relation for any metric, point or probabilistic, as a function of model size or training set size. Scaling studies often find power law dependencies, so it is possible we are already in the saturated regime as we do not see any notable gains. Walmsley et al. (2024) found that while pre-training on ImageNet data can improve downstream galaxy morphology tasks and does scale with data set and model size, most performance gains come from additional pre-training on galaxy images. Investigation into this effect for DeepDISC will be a promising avenue of future work, especially as pre-training using galaxy images has been shown to improve the detection of galactic features through a Faster-RCNN network (Popp et al., 2024). Pre-training on simulated images (such as those used in this study) may help with training the model on real data, which is incomplete in redshift and color-magnitude space due to spectroscopic selection effects. Training sample augmentation has been shown to mitigate this bias and it would be worthwhile to explore whether DeepDISC would benefit from this method. Moskowitz et al. (2024) augmented a photometric catalog to produce a new training set by using a normalizing flow, and such a catalog could be used to simulate new images to produce an augmented training set for DeepDISC.
We have demonstrated that DeepDISC is able to learn more information relevant for photo-z estimation than catalog photometry can provide. This may help in redshift regimes where strong spectral features are not visible through the available filters, although the data used in this study is idealized in some ways compared to real morphological profiles and SEDs. We will continue to test our model on real data for practical implementations. Individual photo-z PDFs are useful for a variety of science cases, but for future applications that use $n(z)$ such as weak lensing, thorough investigation of the selection biases of DeepDISC will be done to properly calibrate measurements. DeepDISC has shown promising performance on simulated LSST data, which we aim to continue to improve and vet as we quickly approach Rubin first light.
@@ Acknowledgements
This paper has undergone an internal review by the LSST DESC, and we thank the internal reviewers, Eric Gawiser and Ismael Mendoza, for thoughtful feedback that improved the paper. We thank Athol Kemball, Matias Carrasco Kind, Yuxiong Wang and Jeff Newman for illuminating discussions, and the anonymous referees for helpful comments. We thank Shirui Luo, Dawei Mu, and Volodymyr Kindratenko at the National Center for Supercomputing Applications (NCSA) for their assistance with the GPU cluster used in this work. G.M., X.L., Y.E. and Y.L. acknowledge support from the NCSA Faculty Fellowship, NCSA Student Pushing INnovation internship program, LSST-DA through grant 2023-SFF-LFI-03-Liu, NSF grant AST-2308174, and NASA grant 80NSSC24K0219. G.M. thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, and the Moore Foundation; his participation in the program has benefited this work. This work was conducted as part of a LINCC Frameworks Incubator. LINCC Frameworks is supported by Schmidt Sciences. AIM, TZ, DO, OL, DB, JK, and MD are supported by Schmidt Sciences. C.J.B. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303803. This material is based upon work supported by the National Science Foundation under Award No. 2303803. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. This work utilizes resources supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign.